An electric motor contains a 250-turn circular coil 6.4 cm in diameter

If it is to develop a maximum torque of 1.2 N⋅m at a current of 3.9 A , what should be the magnetic field strength? (answer in mT)
Answer #1

Here,

let the magnetic field is B

N = 250

area of the coil , A = pi * (0.064/2)^2

I = 3.9 A

maximum torque = N * I * A * B

1.2 = 250 * 3.9 *pi * (0.064/2)^2 * B

solving for B

B = 0.382 T

B = 382 mT

the magnetic field needed is 382 mT